# SFY — A lightweight, high-frequency and phase-resolving wave-buoy for coastal waters

### v1

@@ Abstract
Small lightweight wave buoys, SFYs, designed to operate near the coast, have been developed. The buoys are designed to record and transmit the full time series of surface acceleration at $52$ Hz. The buoy uses the cellular network to transfer data and position (up to $80$ km from the base station). This reduces costs and increases band-width. The low cost and low weight permits the buoys be deployed easily, and in arrays in areas where satellite and wave models struggle to resolve wave and current interaction. The buoys are tested in a wave-flume, the open water and in the breaking waves of the surf. The conditions range from calm to significant wave heights exceeding $7$ m and crashing breakers with accelerations exceeding $10$ g. The high sample rate captures the impulse of breaking waves, and allow them to be studied in detail. Breaking waves are measured and quantified in the open water. We measure breaking waves in the surf and recover the trajectory of waves breaking in the field to a higher degree than previously done. The time series of surface elevation, and accurate positioning, permits adjacent buoys to be correlated in a coherent phase-resolved way. Finally, we offer an explanation and solution for the ubiquitous low-frequency noise in IMU-based buoys and discuss necessary sampling and design to measure in areas of breaking waves.
@@ 1 Introduction
Wave breaking is the primary process by which energy, momentum and gases are exchanged between the air and the sea. It is the limiting factor in wave growth and helps transfer energy through the wave spectrum from shorter to longer waves (Melville, 1996; Babanin, 2009; Deike, 2022). The forces induced by breaking waves are one of the dimensioning factor in marine and nautical engineering, and their effect on coastlines is profound. Most maritime activity happens close to the coast, with more than 60 % of humans living near the coast (Boehm et al., 2017). Estimates further suggest that $80$ of all search and rescue cases are found within $20$ km of the shoreline (Breivik and Allen, 2008). Winds, waves, and currents are heavily modified and influenced by the coastline where waves and currents interact strongly (Halsne et al., 2022, 2023). Consequently, the regions with the most complex wave conditions typically hold the greatest societal importance and are also the most difficult to sample.
Of the many ways to measure waves, in situ heave buoys and wave staffs are the most popular (Holthuijsen, 2010). There exists a range of high–quality commercial moored and drifting wave buoys [see, e.g., the buoys by Datawell (Datawell BV, 2023), SOFAR Ocean (Raghukumar et al., 2019), and Aanderaa (Dorgeville et al., 2018), LainePoiss (Alari et al., 2022)]. These buoys are designed for operational application and therefore have a need to function robustly for long stretches of times. They usually report statistical properties, like significant wave height, and in some cases time series up to a few Hz. Often the time series are only available onboard, and not over telemetry. However, the last decade has seen a small revolution in low-cost, open-source wave-measuring devices (Thomson et al., 2023; Rabault et al., 2020; Feddersen et al., 2023a). These designs generally do not seek to have the same robustness as commercial buoys, but can measure in quite different conditions. Because of the low cost and lower environmental impact, they are also often considered expendable. This is important when measuring in regions where the risk of loss and damage is high. More importantly, these instruments are typically an one or more orders of magnitude smaller and lighter than traditional wave buoys and can thus measure in locations where traditional buoys cannot. Crucially, they also offer the opportunity to measure quantities that larger buoys are incapable of capturing, such as the acceleration in breaking waves (Feddersen et al., 2023b). The limits and capabilities of the instrument determine which questions can be answered. Heterogeneous and different solutions are valuable in themselves in that they have different biases, cover different use cases and make results more robust.
@@
Some open-source buoys have telemetry so that the data can be transmitted to shore. This is typically done using Iridium satellites and the Iridium Short Burst Data (SBD) system, which admit only very low bandwidth (except with high-power, heavier, equipment). Some buoy designs can only store the information on board and thus require the buoy to be retrieved. This is often not possible or requires much time and resources. Loss of data is also an issue with such an approach, especially for deployments in rough conditions. The SFY avoids both these problems by relying on the cellular network to transmit data. This allows us to transmit much more data so that we can send continuous high-frequency time series of the full three-dimensional acceleration vector of the buoy. Additionaly, this allows us to build buoys at a much lower cost and with a smaller physical and environmental footprint. Together this makes it possible to analyse the full time series for impulses incurred by breaking waves, and to resolve the phase of waves between buoys deployed in dense spatial networks. In the coastal environment, resolving this spatial variability is as important as (if not more than) having a single high–quality point measurement. The disadvantage is that the deployments must be within the cellular coverage and will not work far from the coast.
@@
Here, $\omega=2\pi f$ is the angular frequency [rad/s], $f$ is the linear frequency [Hz] and $\mathcal{F}$ is the Fourier transform. When integrating spectra of acceleration density ($A$) to elevation energy (variance density spectrum), $E$, Eq. (2) becomes
@@
There are inherent limitations to measuring a single point of the surface, which cannot capture the changing interference patterns produced by the dispersive wave field (Magnusson et al., 1999). These are better resolved with methods that instantaneously span the spatial field, like stereo-video measurements (see, e.g., Benetazzo (2006); Malila et al. (2022); Malila (2022)) or densely sampled point measurements (e.g Donelan et al. (1985, 1996)). Additionally, the motion of wave buoys tend to underestimate the wave crests (Forristall, 2000); moored buoys are affected by the pull of the anchoring and have a tendency to skirt around the highest wave crests, while drifting buoys overestimate the mean water level by spending too much time near the crests.
@@
$f_{\mathrm{gc}}=\frac{\sqrt{g/k}}{\lambda}=\sqrt{\frac{g}{2\pi\lambda}}.$ (6)
For the free-drifting buoy $f_{\mathrm{gc}}=2.24$ Hz, and for the moored buoy $f_{\mathrm{gc}}=1.09$ Hz. These frequencies will be compared with the spectra from experiments in Sections 3 and 4.
@@ 2 Buoy design
The circular electronics board measures 60 mm in diameter and about 2 cm in height. Figure 1 shows the components with the 800 ml bottle used for housing.
The sensor pipeline generally follows the setup for the OpenMetBuoy (Rabault et al., 2022), hereafter OMB, and benefits from the extensive field testing and validation that the OMB has been subjected to. The inertial measurement unit (IMU) used is either the ISM330DHCX or LSM6DSOX (STMicroelectronics, 2019, 2020), both of which yield gyroscope and accelerometer readings. The IMU low-pass filters the signal internally according to the configured output rate. The theoretical noise level of the ISM330DHCX is typically $60\,\mu g/\mathrm{Hz}^{1/2}$ (maximum $100\,\mu g/\mathrm{Hz}^{1/2}$), according to the manufacturer datasheetSTMicroelectronics (2020). Here, $g=9.81\,\mathrm{m\,s^{-2}}$ is the earth’s gravitational acceleration.
Accelerometer and gyroscope readings from the IMU are fused by applying an Attitude and Heading reference system (AHRS) algorithm implemented with a Kalman filter so that the acceleration is obtained in a North-East-Down (NED) frame of reference (NXP Semiconductors, 2016; Adafruit, 2023). No magnetometer is currently included, so the North-East vector points in an arbitrary direction, but remains stable on short time scales thanks to the measurements provided by the gyroscope. In this analysis only the relative horizontal acceleration is utilized. The IMU is sampled at $208$ Hz and the samples are then fed into the Kalman filter at this same frequency. Finally, the output from the Kalman filter is filtered using a finite-impulse response (FIR) low-pass filter (129 taps with a Hamming window) and decimated down to $52$ Hz or $26$ Hz depending on the application.
It is difficult to quantify the theoretical noise level of the processing chain as a whole since both the accelerometer and gyroscope introduce noise before these are fused in a complex nonlinear way by the Kalman filter. Hence, the noise level of the accelerometer is used as a lower bound of the noise and as an estimate of its frequency dependence. The noise level of the sensors and the processing pipeline as a whole is discussed in more detail by Rabault et al. (2023). Details on how the data are packed, transmitted and received are described in appendix C.
Figure 1: The components that make up the SFY wave buoy. From the left: housing (plastic bottle) with foam for additional stability, battery pack (Alkaline), main board with SD-card, Microcontroller Unit (MCU) in red, modem and GPS in green, IMU, and GPS and cellular antennas.
@@ 3 Lab experiment: Wave flume
Table 1: Buoy and wave gauge configuration for lab experiment.
(a) $f=0.10$ Hz, $A=3.0$ cm (run 6).
(b) $f=0.15$ Hz, $A=5.0$ cm (run 5).
(c) $f=0.25$ Hz, $A=5.0$ cm (run 4).
(d) $f=0.3$ Hz, $A=5.0$ cm (run 3b).
(e) $f=0.5$ Hz, $A=5.0$ cm (run 2).
(f) $f=1$ Hz, $A=5.0$ cm (run 1).
(g) $f=1.5$ Hz, $A=2.0$ cm (run 8a), B16 fixed with rope.
Figure 2: Welch spectrum measured by buoys and wave gauge (WG) measured in a wave flume with piston configured at different frequencies and wave amplitudes. The red vertical line marks frequency where the wavelength is twice the size of the collar width of the buoys ($2.24$ Hz, see Eq. (6)). The vertical dashed lines mark the input frequency and its higher harmonics.
@@
The Welch spectrum in Figure 2 is calculated with a segment length of 4096 for the buoys, and 8192 for the wave gauge which samples at approximately twice the frequency of the wave-buoy.
@@
The measured peaks in the spectra (Figure 2) match the configured piston frequency and its higher harmonics (gray vertical lines) well. The frequency of the wave components measured with the wave gauge also match very well. In the initial experiments the buoys were freely moving at different positions in the wave flume and were subjected to considerable lateral movement and rotation due to Stokes drift, so the phase between the buoys should not be expected to match. Additionally, the vertical amplitude of the generated wave varied considerably laterally along the flume, probably due to interference and reflection (particularly for the lower frequencies).
@@
At approximately $2.2$ Hz there is a marked decrease in the wave-buoy spectra compared to the wave gauge. The collar-width of the buoy (Figure 1) is about $15.5$ cm. From Eq. (6), we see that a wavelength twice that has a deep-water frequency of $2.24$ Hz (red vertical line in Figure 2). The measured amplitude is therefore increasingly dampened as the wavelength approaches the diameter of the buoy. According to Eq. (5), the heave response frequency is $f_{\mathrm{hr}}=3.56$ Hz, but it is very sensitive to the degree of immersion and displacement volume. It seems that the dominating effect on the heave response is the diameter of the buoy, rather than the weight and volume. Yurovsky and Dulov (2020) observe a similar geometric cut-off frequency of about $2.5$ Hz (corresponding to the wavelength twice the hull diameter of their buoy). As they mention, this means that shorter waves can be measured compared to most other alternatives (e.g., $1.2$ Hz for the Spotter buoy (Raghukumar et al., 2019)).
@@
The buoys measure acceleration and have an approximately constant noise level on the measured acceleration. The wave gauge measures displacement and has an approximately constant noise level on the displacement. The difference becomes apparent in the spectrum (Figure 2) at lower frequencies, where the noise level of the wave gauge is flat (horizontal), while the noise level of the wave buoys is proportional to $f^{-2}$ following the double integration from acceleration to displacement. At lower frequencies the wave buoys therefore have a higher noise ratio compared to the wave gauge, whereas at higher frequencies the wave buoy performs better since acceleration is better for measuring high frequency waves (with higher acceleration than displacement), and conversely the wave gauge is better at measuring low-frequency waves. See Sec. 1 and 66.1 for more details.
@@ 3.1 Significant wave height
In Table 2 the significant wave height is calculated from the zeroth moment of the elevation variance spectrum ($H_{m0}$, Eq. 4) and compared to the programmed wave amplitude of the piston, and the significant wave height ($H_{m0}$) of the wave gauge (WG). The theoretical $H_{m0}$ for the generated signal with a single harmonic wave with amplitude $A_{p}$ and variance $\frac{1}{2}A_{p}^{2}$ is calculated as (see Holthuijsen (2010), Chapter 3.5):
@@
The measured signal from the buoys need to be high-pass filtered at a cut-off frequency to avoid low-frequency noise. The wave gauge is not filtered, since it has a flat noise curve on displacement (Figure 2). Automatic selection of low-end cut-off frequency is generally hard. This is a problem that affects IMU based buoys in general (see, e.g., Nose et al. (2018)) and will be discussed further in Section 66.1. Alari et al. (2022) chose the bandpass cut-off frequencies of the measured signal in the wave tank tests to be $0.30$ and $1.28$ Hz and the LainePoiss® buoy was evaluated at frequencies between $0.57$ and $0.73$ Hz.
In Table 2 the cut-off frequency is found automatically by searching for the first spectral minimum above $0.05$ Hz. The algorithm is based on the process outlined in Rabault et al. (2022) (Figure 7) as well as in Tucker (1958); Waseda et al. (2018). However, when this algorithm is unsuccessful it results in unphysically large amplitudes and significant wave height estimates. The auto-detected cut-off frequency is shown and listed in Figure 2. For instance for $1$ Hz in Table 2 and Figure 1(f) the significant wave height for B23 is obviously wrong and overestimated compared to WOB1. This can be attributed to the cut-off frequency being chosen too low for B23, while better for WOB1 (see $f_{c}$ in Figure 1(f), note that the cut-off for B23 is off the scale of the plot). In general it is easier to auto-detect the cut-off frequency when the waves are higher and the signal to noise ratio is greater. Since we know the input frequency for these experiments we can manually choose cut-off frequencies. In Table 3 the significant wave height is shown for the manually chosen frequency band. These values match better, and do not show exaggerated significant wave heights like those present in Table 2 (where a too low or too high cut-off frequency has been picked).

### v2

@@ Abstract
Small lightweight wave buoys, SFYs, designed to operate near the coast, have been developed. The buoys are designed to record and transmit the full time series of surface acceleration at $52$ Hz. The buoy uses the cellular network to transfer data and position (up to $80$ km from the base station). This reduces costs and increases band-width. The low cost and low weight permits the buoys to be deployed easily, and in arrays in areas where satellite and wave models struggle to resolve wave and current interaction. The buoys are tested in a wave-flume, the open water and in the breaking waves of the surf. The conditions range from calm to significant wave heights exceeding $7$ m and crashing breakers with accelerations exceeding $10$ g. The high sample rate captures the impulse of breaking waves, and allows them to be studied in detail. Breaking waves are measured and quantified in the open water. We measure breaking waves in the surf and recover the trajectory of waves breaking in the field to a higher degree than previously done. The time series of surface elevation, and accurate positioning, permits the signal of adjacent buoys to be correlated in a coherent phase-resolved way. Finally, we offer an explanation and solution for the ubiquitous low-frequency noise in IMU-based buoys and discuss necessary sampling and design to measure in areas of breaking waves.
@@ 1 Introduction
Wave breaking is the primary process by which energy, momentum and gases are exchanged between the air and the sea. It is the limiting factor in wave growth and helps transfer energy through the wave spectrum from shorter to longer waves (Melville, 1996; Babanin, 2009; Deike, 2022). The forces induced by breaking waves are one of the dimensioning factors in marine and nautical engineering, and their effect on coastlines is profound. Most maritime activity happens close to the coast, with more than 60 % of humans living near the coast (Boehm et al., 2017). Estimates further suggest that $80$ of all search and rescue cases are found within $20$ km of the shoreline (Breivik and Allen, 2008). Winds, waves, and currents are heavily modified and influenced by the coastline where waves and currents interact strongly (Halsne et al., 2022, 2023; Halsne, 2024). Consequently, the regions with the most complex wave conditions typically hold the greatest societal importance and are also the most difficult to sample.
Of the many ways to measure waves, in situ heave buoys and wave staffs are the most popular (Holthuijsen, 2010). There exists a range of high–quality commercial moored and drifting wave buoys [see, e.g., the buoys by Datawell (Datawell BV, 2023), SOFAR Ocean (Raghukumar et al., 2019), and Aanderaa (Saetre et al., 2023), LainePoiss (Alari et al., 2022)]. These buoys are designed for operational application and therefore have a need to function robustly for long stretches of times. They usually report statistical properties, like significant wave height, and in some cases time series up to a few Hz. Often the time series are only available onboard, and not over telemetry. However, the last decade has seen a small revolution in low-cost, open-source wave-measuring devices (Thomson et al., 2023; Rabault et al., 2020; Feddersen et al., 2023a; Collins et al., 2024). These designs generally do not seek to have the same robustness as commercial buoys, but can measure in quite different conditions. Because of the low cost and lower environmental impact, they are also often considered expendable. This is important when measuring in regions where the risk of loss and damage is high. More importantly, these instruments are typically one or more orders of magnitude smaller and lighter than traditional wave buoys and can thus measure in locations where traditional buoys cannot. Crucially, they also offer the opportunity to measure quantities that larger buoys are incapable of capturing, such as the acceleration in breaking waves (Feddersen et al., 2023b). The limits and capabilities of the instrument determine which questions can be answered. Heterogeneous and different solutions are valuable in themselves in that they have different biases, cover different use cases and make results more robust.
@@
Some open-source buoys have telemetry so that the data can be transmitted to shore. This is typically done using Iridium satellites and the Iridium Short Burst Data (SBD) system, which admit only very low bandwidth (except with high-power, heavier, equipment). Some buoy designs can only store the information on board and thus require the buoy to be retrieved. This is often not possible or requires much time and resources. Loss of data is also an issue with such an approach, especially for deployments in rough conditions. The SFY avoids both these problems by relying on the cellular network to transmit data.
The main improvements of the SFY buoy is that it allows us to transmit much more data, is lightweight, and low cost with a small physical and environmental footprint. This combination provides a number of advantages: It can transmit continuous high-frequency time series of the full three-dimensional acceleration vector measured by the buoy. It can be deployed easily from surf-boards or small vessels, or in lightweight moorings. It can be deployed in arrays of many buoys capturing the spatial variability, which is perhaps more important than a single high–quality point measurement in coastal environments. And it can resolve the phase, impulses and trajectory of waves and breaking waves, also between buoys in the same dense spatial network. The buoy does not currently have a magnetometer, but tracks the orientation in its own reference system on short time scales. The deployments must be within the cellular coverage and will not work far from the coast unless there is a marine network available.
@@
Here, $\omega=2\pi f$ is the angular frequency [rad/s], $f$ is the linear frequency [Hz] and $\mathcal{F}$ is the Fourier transform. When analyzing spectra of elevation energy (variance density spectrum), $E$, no detrending and filtering is necessary. The integration of the acceleration density ($A$), Eq. (2), is performed in one step:
@@
There are inherent limitations to measuring a single point of the surface, which cannot capture the changing interference patterns produced by the dispersive wave field (Magnusson et al., 1999). These are better resolved with methods that instantaneously span the spatial field, like stereo-video measurements (see, e.g., Benetazzo (2006); Malila et al. (2022); Malila (2022)) or densely sampled point measurements (e.g Donelan et al. (1985, 1996)). Additionally, the motion of wave buoys tend to underestimate the wave crests (Longuet-Higgins, 1986; Forristall, 2000); moored buoys are affected by the pull of the anchoring and have a tendency to skirt around the highest wave crests, while drifting buoys overestimate the mean water level by spending too much time near the crests.
@@
$f_{\mathrm{gc}}=\frac{\sqrt{g/k}}{\lambda}=\sqrt{\frac{g}{4\pi r}}.$ (6)
For the free-drifting buoy $f_{\mathrm{gc}}=2.24$ Hz, and for the moored buoy $f_{\mathrm{gc}}=1.09$ Hz. In shallower waters a different phase speed should be used to get a more accurate cut-off frequency. These frequencies will be compared with the spectra from experiments in Sections 3 and 4.
@@ 2 Buoy design
The circular electronics board measures 60 mm in diameter and about 2 cm in height. Figure 1 shows the components with the 800 ml bottle used for housing. Some additional weight is added in order to position the bottle upright in the water. The electronics are placed in bags of bubble wrap, and the antennas positioned as high and unobstructed as possible near the cap. The cap is sealed with self-amalgamating tape. This setup is robust against high impacts and keeps the water out.
The sensor pipeline generally follows the setup for the OpenMetBuoy (Rabault et al., 2022), hereafter OMB, and benefits from the extensive field testing and validation that the OMB has been subjected to. The inertial measurement unit (IMU) used is either the ISM330DHCX or LSM6DSOX (STMicroelectronics, 2019, 2020), both of which yield gyroscope and accelerometer readings. The IMU low-pass filters the signal internally according to the configured output rate. The theoretical noise level of the ISM330DHCX is typically $60\,\mu g/\mathrm{Hz}^{1/2}$ (maximum $100\,\mu g/\mathrm{Hz}^{1/2}$), according to the manufacturers data sheet (STMicroelectronics, 2020). Here, $g=9.81\,\mathrm{m\,s^{-2}}$ is the earth’s gravitational acceleration.
Accelerometer and gyroscope readings from the IMU are fused by applying an Attitude and Heading reference system (AHRS) algorithm implemented with a Kalman filter so that the acceleration is obtained in a North-East-Down (NED) frame of reference (NXP Semiconductors, 2016; Adafruit, 2023). No magnetometer is currently included, so the North-East vector points in an arbitrary direction, but remains stable on short time scales thanks to the measurements provided by the gyroscope. The horizontal direction is therefore only analyzed on short timescales in the buoy reference system (See Sec. 23.2). The IMU is sampled at $208$ Hz and the samples are then fed into the Kalman filter at this same frequency. Finally, the output from the Kalman filter is filtered using a finite-impulse response (FIR) low-pass filter (129 taps with a Hamming window) and decimated down to $52$ Hz or $26$ Hz depending on the application.
It is difficult to quantify the theoretical noise level of the processing chain as a whole since both the accelerometer and gyroscope introduce noise before these are fused in a complex nonlinear way by the Kalman filter. Hence, the noise level of the accelerometer is used as a lower bound of the noise and as an estimate of its frequency dependence. The noise level of the sensors and the processing pipeline as a whole are discussed in more detail by Rabault et al. (2023). Details on how the data are packed, transmitted and received are described in appendix C.
Figure 1: The components that make up the SFY wave buoy. From the left: housing (plastic bottle) with foam for additional stability, battery pack (alkaline), main board shown twice (both sides): Microcontroller Unit (MCU) in red, modem and GPS in green, SD-card, IMU, and GPS and cellular antennas. For scale: the ruler is 60 cm long.
@@ 3 Lab experiment: Wave flume
Table 1: Buoy and wave gauge configuration for lab experiment. See appendix C for more information on IMU frequency discrepancies.
[…]
@@
The Welch spectrum in Fig. 2 is calculated with a segment length of 4096 for the buoys, and 8192 for the wave gauge which samples at approximately twice the frequency of the wave-buoy, both are detrended using a linear least square fit [^2] https://docs.scipy.org/doc/scipy/reference/generated/scipy.signal.welch.html.
@@
The measured peaks in the spectra (Fig. 2) match the configured piston frequency and its higher harmonics (dashed gray vertical lines) well. The frequency of the wave components measured with the wave gauge also match very well. In the initial experiments the buoys were freely moving at different positions in the wave flume and were subjected to considerable lateral movement and rotation due to Stokes drift, so the phase between the buoys should not be expected to match. Additionally, the vertical amplitude of the generated wave varied considerably laterally along the flume, probably due to interference and reflection (particularly for the lower frequencies).
@@
At approximately $2.2$ Hz there is a marked decrease in the wave-buoy spectra compared to the wave gauge. The collar-width of the buoy (Fig. 1) is about $15.5$ cm. From Eq. (6), we see that a wavelength twice that has a deep-water frequency of $2.24$ Hz (red vertical line in Fig. 2). The measured amplitude is therefore increasingly dampened as the wavelength approaches and becomes shorter than the diameter of the buoy. According to Eq. (5), the heave response frequency is $f_{\mathrm{hr}}=3.56$ Hz, but it is very sensitive to the degree of immersion and displacement volume. It seems that the dominating effect on the heave response is the diameter of the buoy, rather than the weight and volume. Yurovsky and Dulov (2020) observe a similar geometric cut-off frequency of about $2.5$ Hz (corresponding to the wavelength twice the hull diameter of their buoy). As they mention, this means that shorter waves can be measured compared to most other alternatives (e.g., $1.2$ Hz for the Spotter buoy (Raghukumar et al., 2019)).
@@
The buoys measure acceleration and have an approximately constant noise level on the measured acceleration. The wave gauge measures displacement and has an approximately constant noise level on the displacement. The difference becomes apparent in the spectrum (Fig. 2) at lower frequencies, where the noise level of the wave gauge is flat (horizontal), while the noise level of the wave buoys is proportional to $f^{-2}$ following the double integration from acceleration to displacement. At lower frequencies the wave buoys therefore have a higher noise ratio compared to the wave gauge, whereas at higher frequencies the wave buoy performs better since acceleration is better for measuring high frequency waves (with higher acceleration than displacement). See Sec. 1 and 66.1 for more details.
@@ 3.1 Significant wave height
In Table 2 the significant wave height is calculated from the zeroth moment of the elevation variance spectrum ($H_{m0}$, Eq. (4)) and compared to the programmed wave amplitude of the piston, and the significant wave height ($H_{m0}$) of the wave gauge (WG). The theoretical $H_{m0}$ for the generated signal with a single harmonic wave with amplitude $A_{p}$ and variance $\frac{1}{2}A_{p}^{2}$ is calculated as (see Holthuijsen (2010), Chapter 3.5):
@@
The measured signal from the buoys need to be high-pass filtered at a cut-off frequency to avoid low-frequency noise. The wave gauge is not filtered, since it has a flat noise curve on displacement (Fig. 2). Automatic selection of low-end cut-off frequency is generally hard. This is a problem that affects IMU based buoys in general (see, e.g., Nose et al. (2018)) and will be discussed further in Section 66.1. Alari et al. (2022) chose the bandpass cut-off frequencies of the measured signal in the wave tank tests to be $0.30$ and $1.28$ Hz and the LainePoiss® buoy was evaluated at frequencies between $0.57$ and $0.73$ Hz.
In Table 2 the cut-off frequency is found automatically by searching for the first spectral minimum above $0.05$ Hz. The algorithm is based on the process outlined in Rabault et al. (2022) (Figure 7) as well as in Tucker (1958); Waseda et al. (2018). However, when this algorithm is unsuccessful it results in unphysically large amplitudes and significant wave height estimates. The auto-detected cut-off frequency is shown and listed in Fig. 2. For instance for $1$ Hz in Table 2 and Fig. 2f the significant wave height for B23 is obviously wrong and overestimated compared to WOB1. This can be attributed to the cut-off frequency being chosen too low for B23, while better for WOB1 (see $f_{c}$ in Fig. 2f, note that the cut-off for B23 is off the scale of the plot). In general it is easier to auto-detect the cut-off frequency when the waves are higher and the signal to noise ratio is greater. Since we know the input frequency for these experiments we can manually choose cut-off frequencies. In Table 3 the significant wave height is shown for the manually chosen frequency band. These values match better, and do not show exaggerated significant wave heights like those present in Table 2 (where a too low or too high cut-off frequency has been picked). In unknown conditions it is therefore best to start with a minimum of $0.05$ Hz for this buoy, and possibly other IMU based buoys. If the wave height suddenly jumps to unrealistic levels or there is significant energy at very low frequencies this is a sign that it should be increased, or that the measurements have been compromised in other ways (e.g. a poor mooring, unstable float or interference from traffic).